# LastingBench: Defend Benchmarks Against Knowledge Leakage

### v1

@@ Appendix B Prompt Templates
Answer the question based on the given passages. The following are the passages: {Context} Answer the question based on the given passages. Question: {Question} Please first provide your answer in the format of Answer:[Your answer]. Then provide your reasoning process step-by-step.(Only include explicit clues) At the end of each reasoning step, include a new line that specifies the key information or reference content used in that step. Please ensure that the [reference content] you include is the complete original sentence or consecutive sentences from the text. Please do not change the punctuation. Do not use ellipses inside the sentence. Follow this format: Answer: [Your answer] Step-by-step Reasoning: 1. [Reasoning step 1] [replaced by your reference content] 2. [Reasoning step 2] [replaced by your reference content]
@@
You are a creative contrarian. Given the question below, and the original answer, first propose a concise alternative answer—that is, a plausible but intentionally misleading answer. Followed are some sentences supporting the original answer, please rewrite them. When rewriting each sentence, modify only the parts necessary to support the antifact answer. Parts unrelated to the answer must keep their original meaning. Be sure that the modified evidence sentences are sufficient to answer the original question. Output must be strictly in the specified JSON format, with no additional text. { "answer": "<your antifact answer here, just provide the answer phrase, no need for complete sentence>", "revised": [ "<rewritten sentence 1>", "<rewritten sentence 2>", ] } Question: {Question} Original answer: {Original Answer} Sentences to rewrite: {Numbered Sentences}

### v2

@@ Appendix B Prompt Templates
Answer the question based on the given passages. The following are the passages: {Context}
Answer the question based on the given passages.
Question: {Question}
Please first provide your answer in the format of Answer:[Your answer]. Then provide your reasoning process step-by-step.(Only include explicit clues)
At the end of each reasoning step, include a new line that specifies the key information or reference content used in that step. Please ensure that the [reference content] you include is the complete original sentence or consecutive sentences from the text. Please do not change the punctuation. Do not use ellipses inside the sentence. Follow this format: Answer: [Your answer] Step-by-step Reasoning: 1. [Reasoning step 1] [replaced by your reference content] 2. [Reasoning step 2] [replaced by your reference content]
@@
You are a creative contrarian. Given the question below, and the original answer, first propose a concise alternative answer—that is, a plausible but intentionally misleading answer. Followed are some sentences supporting the original answer, please rewrite them. When rewriting each sentence, modify only the parts necessary to support the antifact answer. Parts unrelated to the answer must keep their original meaning. Be sure that the modified evidence sentences are sufficient to answer the original question. Output must be strictly in the specified JSON format, with no additional text. { "answer": "<your antifact answer here, just provide the answer phrase, no need for complete sentence>", "revised": [ "<rewritten sentence 1>", "<rewritten sentence 2>", ] } Question: {Question}
Original answer:
{Original Answer}
Sentences to rewrite:
{Numbered Sentences}